OpenStudy (anonymous):

log5(x)-log5(x-2)=3 how do you solve this?

Nnesha (nnesha):

do you know the log properties ??

OpenStudy (anonymous):

im not too sure what they are, my teacher isnt very good at explaining

Nnesha (nnesha):

okay here are log properties quotient rule\[\huge\rm log_b y - \log_b x = \log_b \frac{ x }{ y}\] to condense you can change subtraction to division product rule \[\huge\rm log_b x + \log_b y = \log_b( x \times y )\] addition ----> multiplication power rule \[\huge\rm log_b x^y = y \log_b x\]

Nnesha (nnesha):

look at your question and log properties which one you should apply first ??

OpenStudy (anonymous):

the quotient rule?

Nnesha (nnesha):

yes right bec there is negative sign go ahead and apply that :-)

OpenStudy (anonymous):

log5(x) - log5(x-2) = log5(x/x-2) Then : log5(x/x-2) = 3 which will be : \[\frac{ x }{ x-2 } = 5^{3}\] by simplifying the above equation you will get : x = 125x - 250 then : x = 250/124 Let me know if you got it :)

OpenStudy (anonymous):

ok so i got it up to log5(x-2/x)=3 but i dont know what to do from there

Nnesha (nnesha):

nope remember first number should be at the top